# Ten Years of ZMap

### v2

@@ 2.1. Empirical Analysis
ZMap usage has increased dramatically over the past four years (Figure 1). From January 1 to March 31, 2024, 35.4% of all Internet-wide IPv4 TCP scan packets originated from ZMap. While it is possible that ICMP and UDP scans use other tools, the vast majority of all scan traffic is TCP-based and ZMap still accounts for 33% of all scan packets with only TCP traffic attributed. ZMap scans follow a different distribution of targeted ports than global scans (Figures 2 and 3). For example, ZMap accounts for only 12% of TCP/23 but 69% of TCP/80 and 73% of TCP/8080. In the most extreme case, 99.5% of traffic targeting TCP/8728 (MikroTik router API) globally is from ZMap, driving it to the sixth most scanned port globally. There are also dramatic regional differences in ZMap adoption (Figure 4). For example, while ZMap accounts for more than 66% of scan packets from U.S. hosts, less than 0.5% of Russian scan packets are from ZMap. The outsized ZMap use in the U.S. is largely due to adoption by American security companies (§2.3). In parallel to our work, Griffioen et al. conducted a more in-depth analysis of how scanning has evolved over the past decade (scanning-ten, ); the study found that 59% of Internet-wide scans in 2024 used ZMap.
@@ 2.2. Academic Research
While ZMap is a general measurement tool, it has most prominently been used by the security community. Notably, ZMap has been used in 38 papers to uncover protocol weaknesses in TLS and underlying cryptographic primitives (beurdouche2015messy, ; adrian2015imperfect, ; aviram2016drown, ) as well as to uncover deployment challenges and measure adoption (nemec2017measuring, ; kranch2015upgrading, ; holz2015tls, ). Collecting X.509 certificates through scans, 28 papers have shed light on the WebPKI prior to the widespread adoption of Certificate Transparency (vandersloot2016towards, ). There is also a large contingent of papers that have measured the Internet exposure of IoT devices (25 papers), ICS (14 papers), and security-relevant services (12 papers). Beyond understanding deployment patterns, a number of papers have been able to identify real-world attacks (durumeric2015neither, ), to reverse engineering attacker infrastructure (antonakakis2017understanding, ; marczak2015pay, ; marczak2014governments, ), and to conduct large-scale security notifications using ZMap data (li2016exploring, ; durumeric2014matter, ). We encourage the security community to embrace research that builds Internet measurement tools and techniques since these are frequently used to understand and improve security practices.
More networking-focused papers cover topics like DNS (24 papers), BGP/RPKI (12 papers), censorship (14 papers), and IP usage and NAT (10 papers). In addition to these studies, 53 other papers reference following the recommended practices when conducting other measurements. We provide more details in Appendix B.
@@ 2.4. Malicious Use
Most security tools have the potential for both helping defenders and being misused by attackers. ZMap is no exception and there is evidence that ZMap has been used maliciously. While it is difficult to ascertain the intent of network traffic from a shared cloud provider without any application-layer visibility (hiesgen2022spoki, ; izhikevich2023cloud, ), past darknet analysis has shown that attackers have used “bulletproof” hosting providers to carry out scans for vulnerable services, including MSSQL, RDP, and Mikrotik’s router API (durumeric2014internet, ; anand2023aggressive, ). Anecdotally, attackers also have begun to repurpose ZMap to carry out DOS attacks (russia-dos, ), and, recently, two IoT botnets incorporated ZMap into their malware: in 2021, Palo Alto Networks found a Mirai variant adopted ZMap to search for SSH services (mirai-zmap, ), and, in 2023, a Medusa variant adopted ZMap to find Telnet services (medusa-zmap, ).
@@ 6. Internet Citizenship
    Investigate whether existing datasets (e.g., from Censys or Rapid7) suffice. Often, these datasets provide better coverage than one-off scans and reduce aggregate bandwidth.
@@
    Deploy WHOIS entries for research subnets that help operators to directly contact researchers.
@@
    Validate how handshakes will appear in logs. For example, some benign SSH handshakes inadvertently show up as failed authentication attempts that conern operators.

### v3

@@ Abstract.
Internet Measurement, Internet Scanning, ZMap
ccs: General and reference Measurementccs: Networks Network protocolsccs: Networks Network monitoringccs: Software and its engineering Software creation and managementccs: Security and privacy Network securityjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM Internet Measurement Conference; November 4–6, 2024; Madrid, Spainbooktitle: Proceedings of the 2024 ACM Internet Measurement Conference (IMC ’24), November 4–6, 2024, Madrid, Spaindoi: 10.1145/3646547.3689012isbn: 979-8-4007-0592-2/24/11
@@ 2.1. Empirical Analysis
ZMap usage has increased dramatically over the past four years (Figure 1). From January 1 to March 31, 2024, 35.4% of all Internet-wide IPv4 TCP scan packets originated from ZMap. While it is possible that ICMP and UDP scans use other tools, the vast majority of scan traffic is TCP-based and ZMap still accounts for 33% of all scan packets with only TCP traffic attributed. ZMap scans follow a different distribution of targeted ports than other scans (Figures 2 and 3). For example, ZMap accounts for only 12% of TCP/23 but 69% of TCP/80 and 73% of TCP/8080. In the most extreme case, 99.5% of traffic targeting TCP/8728 (MikroTik router API) is from ZMap, driving it to the sixth most scanned port. There are also dramatic regional differences (Figure 4). For example, while ZMap accounts for more than 66% of scan packets from U.S. hosts, less than 0.5% of Russian scan packets are from ZMap. The outsized U.S. use is driven by its adoption by American security companies (§2.3). In parallel to our work, Griffioen et al. conducted an in-depth analysis of how scanning has evolved over the past decade (scanning-ten, ); the study found that 59% of Internet-wide scans in 2024 used ZMap.
@@ 2.2. Academic Research
While ZMap is a general measurement tool, it has most prominently been used by the security community (Appendix B). Notably, ZMap has been used in 38 papers to uncover protocol weaknesses in TLS and underlying cryptographic primitives (beurdouche2015messy, ; adrian2015imperfect, ; aviram2016drown, ), and to uncover deployment challenges and measure adoption (nemec2017measuring, ; kranch2015upgrading, ; holz2015tls, ). Collecting X.509 certificates, 28 papers have shed light on the WebPKI prior to the widespread adoption of Certificate Transparency (vandersloot2016towards, ). There is also a large contingent of papers that have measured the exposure of IoT devices (25 papers), ICS (14 papers), and security-relevant services (12 papers). Beyond understanding deployment patterns, a number of papers have been able to identify real-world attacks (durumeric2015neither, ), to reverse engineering attacker infrastructure (antonakakis2017understanding, ; marczak2015pay, ; marczak2014governments, ), and to conduct large-scale notifications using ZMap (li2016exploring, ; durumeric2014matter, ). We encourage the security community to embrace research that builds Internet measurement tools and techniques since these are frequently used to understand and improve security.
Networking-focused papers cover topics like DNS (24 papers), BGP/RPKI (12 papers), censorship (14 papers), and IP usage and NAT (10 papers). In addition to these studies, 53 other papers reference the recommended practices when conducting measurements.
@@ 2.4. Malicious Use
Most security tools have the potential for both helping defenders and being misused by attackers. ZMap is no exception and there is evidence that ZMap has been used maliciously. While it is difficult to ascertain the intent of network traffic from shared providers without application-layer visibility (hiesgen2022spoki, ; izhikevich2023cloud, ), past darknet analysis has shown that attackers have used “bulletproof” hosting providers to carry out scans for vulnerable services, including MSSQL, RDP, and Mikrotik’s router API (durumeric2014internet, ; anand2023aggressive, ). Anecdotally, attackers have repurposed ZMap to carry out DOS attacks (russia-dos, ), and, recently, two IoT botnets incorporated ZMap into their malware: between 2021–2023, variants of the Mirai and Medusa botnet adopted ZMap (mirai-zmap, ; medusa-zmap, ).
@@ 6. Internet Citizenship
    Investigate whether existing datasets suffice. Often, these provide better coverage and reduce aggregate bandwidth.
@@
    Deploy WHOIS entries that identify how to contact you.
@@
    Validate how handshakes will appear in logs. For example, some benign SSH handshakes inadvertently show up as failed authentication attempts that concern operators.